A progressive wave equation is represented by Y=a if the face difference of a progressive wave is 45 degree. Determine the value of x in the equation
Y=A sin (150πt - πx/wavelength)
1
Expert's answer
2020-12-14T12:18:08-0500
"Y=A \sin (150πt - 0.25πx)\\=A \sin \left(\frac{2vt}{\lambda} - \frac{2x}{\lambda}\right)"

The phase difference


"\frac{x}{4}=\frac{2x}{\lambda}\\\lambda=8\ cm\\\frac{2x}{8}=45\degree=\frac{\pi}{4}\\x=\pi\ cm"